Waiting period. Jamal is waiting to be a millionaire. He wants to know how long he must wait if a. he invests $26,065.17 at 18% today? b. he invests $53,570.94 at 13% today? c. he invests $137,127.56 at 11% today? d. he invests $300,240.64 at 3% today? a. How long will Jamal have to wait to become a millionaire if he invests $26,065.17 at 18% today? years (Round to the nearest whole number.)
Answer #1

a. n = ln($1,000,000/$26,065.17) / ln (1.18) = 22 years

b. n = ln($1,000,000/$53,570.94) / ln (1.13) = 24 years

c. n = ln($1,000,000/$137,127.56) / ln (1.11) = 19 years

d. n = ln($1,000,000/$300,240.64) / ln (1.03) = 41 years